OpenStudy (anonymous):

Solve: 3 and one half a=490 or 31/2a=490

OpenStudy (anonymous):

3 and 1/2 = 7/2 so 7/2 *a=490 a=490 * 2/7 a=2*70 a=140

OpenStudy (gw2011):

(31/2)(a) = 490 a = (490)(2/31) a = 980/31 Substituting a=980/31 into (31/2)(a)=490 to verify the correctness: (31/2)(980/31) = 980/2 = 490